# Highly Efficient Self-Adaptive Reward Shaping for Reinforcement Learning

### v1

@@ 4.2 Highly Efficient Beta Distribution Derivation
In this section, we present the methods to obtain the success and failure counts, $N_{S}(s_{i})$ and $N_{F}(s_{i})$, which are required for deriving the Beta distributions. In tabular environments with low-dimensional, discrete, and finite state spaces, tracking these counts can be straightforwardly accomplished using two tables. However, most real-world environments involve high-dimensional, continuous, and infinite state spaces, making it impractical to maintain exact counts for every state.
@@
where $K(\cdot)$ is the kernel function, and $|\mathcal{D}_{X}|$ is the total number of states in the corresponding buffer. We select the Gaussian kernel in our implementation. The estimated density $\tilde{d}_{X}(s_{i})$ serves as a proxy for the likelihood of encountering state $s_{i}$ under success or failure scenarios, providing a statistically sound basis for estimating $N_{X}(s_{i})$. By multiplying $\tilde{d}_{X}(s_{i})$ with the total number of observed states $N$, and representing states as $k$-dimensional vectors, the count $\tilde{N}_{X}(s_{i})$ is estimated as:
$\tilde{N}_{X}(\bm{s}_{i})=N\times\tilde{d}_{X}(\bm{s}_{i})=\frac{N}{|\mathcal{D}_{X}|}\sum_{j=1}^{|\mathcal{D}_{X}|}{\exp\left(-\frac{||\bm{s}_{i}-\bm{s}_{j}||^{2}}{2\sigma^{2}}\right)},\quad X\in\{S,F\},$ (7)
where $\sigma$ is a hyperparameter that controls the bandwidth of the Gaussian kernel.
We further integrate Random Fourier Features (RFF) (Rahimi and Recht, 2007) to reduce computational complexity, as calculating the Gaussian kernel can be expensive, especially in scenarios involving high-dimensional state spaces and large buffers. RFF approximates the kernel function of the original $k$-dimensional states through an inner product of $M$-dimensional randomized features:
@@
where $\bm{I}_{k}$ is the $k\times k$ identity matrix. Equation 9 is applied for the Gaussian kernel, while different kernels, such as Laplacian or Cauchy kernels, have different spectral distributions. Detailed derivations of the RFF method and the list of spectral distributions are provided in Appendix A.
Suppose for each iteration, the buffer size of $\mathcal{D}_{X}$ is $D$ and the batch size is $B$, the computational complexity to retrieve the corresponding count $N_{X}$ is $O(MDB)$, representing a linear complexity (detailed derivation in Appendix B). RFF transforms the nonlinear kernel computations into vectors’ linear operations, significantly accelerating computation by leveraging the vectorization capabilities of GPUs (Dongarra et al., 2014). This demonstrates its higher efficiency compared to other methods that rely on network updates and inferences involving extensive nonlinear computations.
Moreover, we introduce a hyperparameter, the retention rate $\phi\in(0,1]$, to regulate the volume and diversity of states stored in the buffers. Instead of storing all the states in each trajectory, we uniformly retain a specific portion $\phi$. This strategy offers two main benefits: (1) as adjacent states in one trajectory tend to be highly similar, especially those closer to the initial state are repetitive and uninformative, retaining a portion of states can skip redundant states and increase sample diversity; (2) maintaining a lower retention rate in the early training stage keeps lower $N_{S}$ and $N_{F}$, resulting in a broader Beta distribution and avoiding premature overconfidence.
Given the retention rate $\phi$, the space complexity of maintaining two buffers is $O(\phi T|s|)$, where $T$ is the total number of iterations and $|s|$ is the size of a single state. To address scalability concerns in more complex or high-dimensional scenarios, we propose leveraging the existing replay buffer of the backbone algorithm, soft actor-critic. By augmenting this buffer with a flag array that indicates whether each state is a success or failure, while this introduces additional indexing computations during processing, we can efficiently save space requirements.
@@ 4.3 The SASR Mechanism for RL agents
Building upon the SASR reward, we employ the soft actor-critic (SAC) algorithm of Haarnoja et al. (2018a) as the foundation for our RL agent. Let $\pi^{\mathcal{A}_{P}}_{\theta}$ be parameterized policy network and $Q^{\mathcal{A}_{P}}_{\phi}$ be parameterized Q-network. We optimize the Q-function by minimizing the mean squared error (MSE) between the predicted Q-value and the temporal difference (TD) target:
$\mathcal{L}(\phi)=\Big{(}Q_{\phi}(s_{t},a_{t})-\big{(}r^{E}_{t}+\lambda R^{S}(s_{t})+\gamma Q_{{\phi}^{\prime}}(s_{t+1},a_{t+1})\big{)}\Big{)}^{2},$ (10)
where $Q_{{\phi}^{\prime}}$ is obtained from a secondary frozen target network to maintain a fixed objective (Mnih et al., 2015). It is worth noting that the environmental reward $r^{E}_{t}$ is retrieved from the replay buffer. Conversely, the shaped reward $R^{S}(s_{t})$ is computed in real-time using the most recently updated $N_{S}(s_{t})$ and $N_{F}(s_{t})$, thereby capturing the latest advancements in learning progress.
We optimize the policy network by maximizing the expected Q-value and the entropy of the policy $\mathcal{H}\big{(}\pi^{\mathcal{A}_{P}}(\cdot|s_{t})\big{)}$ (Haarnoja et al., 2018b):
$\mathcal{L}(\theta)=\mathbb{E}_{a_{t}\sim\pi_{\theta}(\cdot|s_{t})}\big{[}-Q_{\phi}(s_{t},a_{t})+\log{\pi_{\theta}(a_{t}|s_{t})}\big{]}.$ (11)
The outline of the SASR is summarized in Algorithm 1. We maintain two buffers to store successful and failed states and estimate the success and failure counts from them. Subsequently, shaped rewards are computed by sampling from Beta distributions followed by a scale transformation. The shaped rewards, combined with environmental rewards, are used to update the agent’s Q-function and policy.
In summary, SASR offers two primary advantages: First, it leverages the adaptive properties of Beta distributions to continuously adjust the agent’s confidence levels in understanding the environment, effectively maintaining the balance between exploration and exploitation. Second, SASR operates as a non-parametric, learning-free approach, enabling highly efficient reward calculation.
@@ 5 Experiments
We evaluate SASR on continuous control environments, including seven MuJoCo tasks (Todorov et al., 2012): AntStand, AntSpeed, AntFar, AntVerFar, WalkerKeep, HumanStand, and HumanKeep, two robotic tasks (de Lazcano et al., 2023): RobotReach and RobotPush, as well as one physical simulation task MountainCar (Towers et al., 2023), shown in Figure 2. Our environment setup follows the guidelines established in the ReLara paper (Ma et al., 2024), with ReLara also included as a baseline for comparison. The tasks provide only sparse rewards, with a reward of $1$ granted upon reaching the final objective within the maximum permitted steps, and 00 for all other states. To ensure robust experimental validation, we train multiple instances for each setting using different random seeds and report the average results. Additionally, we maintain uniform hyperparameters and network architectures across all experiments, as detailed in Appendix D.
@@
Figure 2: MuJoCo, robotic and physical simulation tasks with their objectives in our experiments.
@@ 5.1 Comparison and Discussion
Baselines. We compare SASR with eight baselines to benchmark its performance: (a) RL with an Assistant Reward Agent (ReLara) (Ma et al., 2024), (b) RL Optimizing Shaping Algorithm (ROSA) (Mguni et al., 2023), (c) Exploration-Guided RS (ExploRS) (Devidze et al., 2022), (d) Count-based static hashing exploration (#Explo) (Tang et al., 2017), (e) Random Network Distillation (RND) (Burda et al., 2018), (f) Soft Actor-Critic (SAC) (Haarnoja et al., 2018a), (g) Twin Delayed Deep Deterministic Policy Gradient (TD3) (Fujimoto et al., 2018), and (h) Proximal Policy Optimization (PPO) (Schulman et al., 2017). Algorithms (a) to (e) are all reward shaping methods, involving either exploration bonuses or auxiliary agents to provide additional rewards to tackle the sparse-reward challenge, while algorithms (f) to (h) are advanced RL algorithms. For the implementation of these baselines, we employ either the CleanRL library (Huang et al., 2022) or codes provided by the authors.
Figure 3 and Table 1 report the average episodic returns and standard errors throughout training. Our findings indicate that SASR surpasses the baselines in terms of sample efficiency, learning stability, and convergence speed. The primary challenges in these environments are attributed to delayed and sparse rewards alongside long horizons, making exploration essential for timely obtaining successful trajectories. Although the exploration strategies of algorithms such as ExploRS, #Explo, and RND are designed to reward novel states, effectively expanding the early exploration space with the direct additional target, they continue focusing on discovering novel states, overlooking the implicit values of these states, which makes them fail to return to the final objectives.
@@
SASR, on the other hand, utilizes the success rate as auxiliary information to assess each state’s contribution to task completion. This metric is consistent with the agent’s primary objectives, and incorporating this meaningful signal into the temporal difference updating at each step significantly enhances learning efficiency. Without this integration, the agent would have to wait until the end of the trajectory to obtain informative signals. Furthermore, SASR regulates an efficient exploration strategy by injecting substantial random rewards during the early phases to optimize the agent in multiple directions thereby increasing the likelihood of collecting positive samples. This is achieved by sampling from initially wider Beta distributions. Importantly, these Beta distributions adaptively yield increasingly confident and accurate success rates as more data is collected, thereby accelerating the agent’s learning. These strategies collectively enhance SASR’s efficiency and stability in handling these challenging tasks.
While ReLara employed a similar exploration mechanism with SASR by rewards perturbations to encourage exploration, it relies on a separate black-box agent, which requires more iterations to converge. In contrast, SASR’s success rate sampling is more intuitive and direct, reducing the delay in providing valuable information to the policy agent. ReLara’s advantage lies in its integration of the policy agent’s action into the reward construction, as evidenced in the RobotReach task. The target point for the robot is randomly selected in each episode, this can be considered as a higher dimension for SASR, because of the changing target. In this case, ReLara achieves better results, with its access to action information. This can be improved in SASR by incorporating the action as an additional feature in the state vector.
@@ 5.2 Effect of Self-Adaptive Success Rate Sampling
SASR introduces a novel self-adaptive mechanism that balances exploration and exploitation by maintaining the randomness of the shaped rewards. Figure 4 illustrates the generated rewards at five stages of learning in the AntStand task. Figure 4 (bottom) represents the learning curve, and Figure 4 (top) illustrates the distribution of the generated rewards over the "height of the ant" feature, one of the dimensions in the state vector.
Figure 4: Distribution of the shaped rewards over the height of the ant robot in the AntStand task, at different stages of the training process.
As learning progresses, the evolving shaped rewards exhibit two crucial attributes: the values transition from random to meaningful, and the variance decreases from uncertain to deterministic. In the early stages of training, the shaped rewards involve significant randomness due to higher uncertainty. Although these random signals contain limited information, they prompt the agent to take tiny optimization steps in diverse directions, effectively shifting the anchors of the policies to various regions. This strategy broadens the actions generated by SAC’s stochastic policy, encouraging the agent’s exploration and increasing the diversity of collected samples. From another perspective, the perturbations on the rewards can be seen as a relaxation of the assumption that only novel states should be rewarded, instead, randomly reward any state, which also shows its empirical effectiveness. In the later stages, the rewards are observed to converge to closely match the height of the robot, a meaningful and intuition-aligned metric, which enhances the agent’s exploitation of its policy. This self-adaptive mechanism for balancing exploration and exploitation improves SASR’s ability to gather positive samples earlier, thereby facilitating a faster convergence to the optimal policy.
@@ 5.3 Ablation Study
We conduct ablation studies to investigate five key components of SASR: (1) sampling from Beta distributions versus using the deterministic success rates; (2) the retention rate $\phi$, which controls the volume of historical information; (3) the bandwidth $h$ of the kernel density estimation; (4) the number of dimensions $M$ of RFF features; and (5) the weight factor $\lambda$, which adjusts the scales of the shaped rewards. We select six representative tasks to report the experimental results, and the quantitative data is provided in Appendix C.
Sampling from Beta distributions. We examine a variant of SASR that omits sampling from the Beta distribution and instead directly uses the success rate $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$. Comparative results are shown in Figure 5. In the early stages of learning, limited experience makes this success rate an inaccurate estimate of the true value. Using this overly confident fixed value can mislead the agent. Additionally, removing the process of sampling from the Beta distribution disrupts the agent’s exploration encouraged by random rewards, leading to slower convergence. The results highlight the importance of sampling from the Beta distribution for effective learning.
Figure 5: The impact of omitting Beta distribution sampling on the performance of SASR.
Figure 6: Learning performance of SASR on different retention rates $\phi$.
Retention rate $\phi$. The retention rate controls the counts’ estimation, subsequently impacting the Beta distribution and sampled rewards. We assess various retention rates in Figure 6. A high retention rate, such as $\phi=1$, preserving all samples, results in a densely populated and redundant experience pool, causing the Beta distribution to be prematurely overconfident, which can degrade performance. Conversely, a lower retention rate $\phi=0.01$ requires more iterations to accumulate sufficient samples, thereby slowing the convergence to meaningful rewards and affecting the agent’s convergence speed. The results suggest that carefully balancing the retention rate is crucial, and promising future work could involve adaptive adjustment of the retention rate during the learning process.
Bandwidth $h$ of Gaussian kernel. We evaluate different bandwidths $h$ of the Gaussian kernel for density estimation, as shown in Figure 7. Beyond fixed bandwidths, we also test a linearly decreasing bandwidth configuration ($h:0.5\rightarrow 0.1$), which represents progressively increasing confidence in KDE. Results indicate that an optimal bandwidth selection is crucial. A small bandwidth $h=0.01$ results in farther distances between samples, leading many to have zero estimated density; while a large bandwidth $h=1$ causes the samples indistinguishable due to the overly flat kernel function. Both scenarios lead to suboptimal performance. The decreasing bandwidth setting does not yield significant improvement and tends to reduce stability due to inconsistent density estimations.
@@
Figure 7: Learning performance of SASR on different bandwidths $h$ of Gaussian kernel.
Figure 8: Learning performance of SASR on different RFF feature dimensions $M$.
@@
Figure 9: Learning performance of SASR on different shaped reward weight factors $\lambda$.
@@
RFF feature dimensions $M$. Learning performance on different RFF feature dimensions $M$ is shown in Figure 8. SASR shows relatively low sensitivity on $M$, provided it is large enough to capture the complexity of the states. Results indicate that values like $M=500,1000,2000$ all yield comparable performance levels. However, a significantly lower dimension, such as $M=50$, degrades performance, which is also suggested in the original RFF study (Rahimi and Recht, 2007).
@@
Shaped reward weight factor $\lambda$. We evaluate the impact of different shaped reward scales on learning performance, as shown in Figure 9. We observe that SASR performs better with intermediate values like $\lambda=0.4,0.6,0.8$. At $\lambda=0.2$, the minimal shaped reward scale reduces state differentiation, leading to suboptimal performance. At $\lambda=1$, aligning the shaped reward scale with the environmental reward introduces excessive uncertainty and randomness, which may overwhelm the feedback disturbances and hinder learning. The findings highlight that maintaining a balanced reward scale is important for achieving optimal learning outcomes.

### v3

@@ 4.2 Highly Efficient Beta Distribution Derivation
In this section, we present the methods to obtain the success and failure counts, $N_{S}(s_{i})$ and $N_{F}(s_{i})$, for deriving the Beta distributions. In tabular environments with low-dimensional, discrete, and finite state spaces, tracking these counts can be straightforwardly accomplished using two tables. However, most real-world environments involve high-dimensional, continuous, and infinite state spaces, making it impractical to maintain exact counts for every state.
@@
where $K(\cdot)$ is the kernel function and $|\mathcal{D}_{X}|$ is the buffer size. We select Gaussian kernel in our implementation. The estimated density $\tilde{d}_{X}(s_{i})$ approximates the likelihood of encountering state $s_{i}$ in success or failure scenarios, providing a statistically sound basis for estimating $N_{X}(s_{i})$. By multiplying $\tilde{d}_{X}(s_{i})$ by the total number of observed states $N$, the count $\tilde{N}_{X}(s_{i})$ is estimated as:
$\tilde{N}_{X}(\bm{s}_{i})=N\times\tilde{d}_{X}(\bm{s}_{i})=\frac{N}{|\mathcal{D}_{X}|}\sum_{j=1}^{|\mathcal{D}_{X}|}{\exp\left(-\frac{||\bm{s}_{i}-\bm{s}_{j}||^{2}}{2h^{2}}\right)},\quad X\in\{S,F\},$ (7)
where $h$ is a hyperparameter, the bandwidth of the Gaussian kernel.
We further integrate Random Fourier Features (RFF) (Rahimi and Recht,, 2007) to reduce computational complexity, as calculating the Gaussian kernel can be expensive, especially in scenarios involving high-dimensional state spaces and large buffers. RFF approximates the kernel function of the original $k$-dimensional states through an inner product of $M$-dimensional randomized features:
@@
where $\bm{I}_{k}$ is the $k\times k$ identity matrix. Equation 9 is applied for the Gaussian kernel, while different kernels and the detailed derivations of the RFF method are provided in Appendix A.
4.2.1 Time and Space Complexity of SASR
Suppose the buffer size of $\mathcal{D}_{X}$ is $D$ and the batch size is $B$ per iteration, the computational complexity to compute the count $N_{X}$ is $O(MDB)$, indicating linear complexity (detailed derivation in Appendix B). RFF transforms nonlinear kernel computations into linear vector operations, significantly speeding up computation by leveraging the vectorization capabilities of GPUs (Dongarra et al.,, 2014). This demonstrates its higher efficiency compared to methods that rely on network updates and inferences involving extensive nonlinear computations. For an empirical comparison of inference time with and without RFF, as well as related discussions, please refer to Appendix C.
Given the retention rate $\phi$, the space complexity of maintaining two buffers is $O(\phi T|s|)$, where $T$ is the total iterations and $|s|$ is the size of a single state. However, storage space can be largely saved by leveraging the existing replay buffer in off-policy backbone RL algorithms, such as SAC (Haarnoja et al., 2018a, ) and TD3 (Fujimoto et al.,, 2018). By augmenting the replay buffer with a flag that marks each state as either success or failure, we can efficiently manage space requirements despite some additional indexing computations introduced during processing.
4.2.2 Implementation Details
Retention Rate. We introduce a hyperparameter, the retention rate $\phi\in(0,1]$, to regulate the volume and diversity of states stored in the buffers. Rather than storing all encountered states, we uniformly retain a specific portion of $\phi$. The motivation behind this are: (1) adjacent states in one trajectory tend to be highly similar, especially those near the initial state are repetitive and uninformative, retaining a portion of states can skip redundant states and increase sample diversity; (2) using a lower retention rate in the early stage keeps $N_{S}$ and $N_{F}$ lower, resulting in broader Beta distributions and preventing premature overconfidence.
Defining Success and Failure. In tasks where sparse rewards are only given at the end of an episode to indicate task completion, the entire trajectory can be classified as either a success or failure based on the episodic reward. For tasks with sparse rewards that do not explicitly indicate task completion, we segment the trajectories by positive reward occurrences. Specifically, if a reward is obtained within a pre-defined maximum steps, the corresponding sub-sequence is classified as a success; otherwise, it is considered a failure.
@@ 4.3 The SASR Mechanism for RL agents
Building upon the SASR reward, we employ the soft actor-critic (SAC) algorithm by Haarnoja et al., 2018a as the foundation for our RL agent. Let $Q^{\mathcal{A}_{P}}_{\psi}$ be parameterized Q-network and $\pi^{\mathcal{A}_{P}}_{\theta}$ be parameterized policy network. The Q-network can be optimized by the following loss function:
$\mathcal{L}(\psi)=\Big{(}Q_{\psi}(s_{t},a_{t})-\big{(}r^{E}_{t}+\lambda R^{S}(s_{t})+\gamma Q_{{\psi}^{\prime}}(s_{t+1},a_{t+1})\big{)}\Big{)}^{2},$ (10)
where $Q_{{\psi}^{\prime}}$ is obtained from a secondary frozen target network to maintain a fixed objective (Mnih et al.,, 2015). It is worth noting that the environmental reward $r^{E}_{t}$ is retrieved from the replay buffer. Conversely, the shaped reward $R^{S}(s_{t})$ is computed in real-time using the most recently updated $N_{S}(s_{t})$ and $N_{F}(s_{t})$, thereby capturing the latest advancements in learning progress.
We optimize the policy network by maximizing the expected Q-value and the entropy of the policy $\mathcal{H}\big{(}\pi^{\mathcal{A}_{P}}(\cdot|s_{t})\big{)}$, following Haarnoja et al., 2018b :
$\mathcal{L}(\theta)=\mathbb{E}_{a_{t}\sim\pi_{\theta}(\cdot|s_{t})}\big{[}-Q_{\psi}(s_{t},a_{t})+\log{\pi_{\theta}(a_{t}|s_{t})}\big{]}.$ (11)
@@
The flow of the SAC-embedded SASR algorithm is summarized in Algorithm 1. We maintain two buffers to store successful and failed states and estimate the success and failure counts from them. Subsequently, shaped rewards are computed by sampling from Beta distributions followed by a scale transformation. The shaped rewards, combined with environmental rewards, are used to update the agent’s Q-function and policy.
@@ 5 Experiments
We evaluate SASR in continuous control environments, following the guidelines established by Ma et al., 2024b , including seven MuJoCo tasks (Todorov et al.,, 2012), two robotic tasks (de Lazcano et al.,, 2023), and one physical simulation task (Towers et al.,, 2023), as shown in Figure 2. All tasks provide extremely sparse rewards, with a reward of $1$ granted only upon reaching the final objective within the maximum permitted steps. To ensure robust validation, we run $10$ instances for each setting using different random seeds and report the average results. We also maintain consistent hyperparameters and network architectures across all experiments, as detailed in Appendix E.
@@
Figure 2: MuJoCo, robotic and physical simulation tasks with their objectives in our experiments. Detailed descriptions of each task are provided in Appendix G.
@@ 5.1 Comparison and Discussion
Baselines. We compare SASR with eight baselines to benchmark its performance: (a) RL with an Assistant Reward Agent (ReLara) (Ma et al., 2024b, ), (b) RL Optimizing Shaping Algorithm (ROSA) (Mguni et al.,, 2023), (c) Exploration-Guided RS (ExploRS) (Devidze et al.,, 2022), (d) Count-based static hashing exploration (#Explo) (Tang et al.,, 2017), (e) Random Network Distillation (RND) (Burda et al.,, 2018), (f) Soft Actor-Critic (SAC) (Haarnoja et al., 2018a, ), (g) Twin Delayed DDPG (TD3) (Fujimoto et al.,, 2018), and (h) Proximal Policy Optimization (PPO) (Schulman et al.,, 2017). Algorithms (a) to (e) are all reward shaping methods, involving either exploration bonuses or auxiliary agents to shape rewards, while algorithms (f) to (h) are advanced RL algorithms.
Figure 3 and Table 1 report the average episodic returns and standard errors throughout training. Our findings indicate that SASR surpasses the baselines in terms of sample efficiency, learning stability, and convergence speed. The primary challenge in these environments is the extremely sparse reward given after a long horizon, making exploration essential for timely obtaining successful trajectories. Although the exploration strategies of algorithms such as ExploRS, #Explo, and RND are designed to reward novel states, effectively expanding the early exploration space with the direct additional target, they continue focusing on discovering novel states, overlooking the implicit values of these states, which makes them fail to return to the final objectives.
@@
SASR outperforms the baselines primarily due to its self-adaptive reward evolution mechanism. In the early phases, SASR encourages exploration by injecting substantial random rewards to optimize the agent in multiple directions, thereby increasing the likelihood of collecting positive samples. Unlike novelty-driven exploration strategies that reward only novel states, SASR relaxes this assumption by rewarding random states, under the belief that all states are worth exploring, and this approach has proven empirically effective. Additionally, SASR leverages the success rate as auxiliary information to assess each state’s contribution to task completion, aligning with the agent’s primary objectives. As more data is collected, the success rate becomes more accurate, and the shaped reward provides more valuable guidance, significantly enhancing the exploitation and stabilizing the convergence toward optimal policies. Cooperatively, these strategies improve SASR’s sample efficiency and convergence stability in challenging tasks.
While ReLara used a similar exploration mechanism to SASR by perturbing reward functions, it relies on an independent black-box agent, requiring more iterations to converge. In contrast, SASR’s success rate sampling is more direct, reducing the delay in converging to valuable information. ReLara’s advantage lies in incorporating the policy agent’s actions into reward construction, as seen in the RobotReach task, where the target point is randomly selected in each episode. In this case, ReLara outperforms SASR due to access to action information. This can also be achieved in SASR by treating actions as additional features in the state vector.
@@ 5.2 Effect of Self-Adaptive Success Rate Sampling
SASR introduces a novel self-adaptive mechanism that balances exploration and exploitation by maintaining the randomness of the shaped rewards. To further explore the effect of this mechanism, we use the AntStand task as a case study, analyzing the shaped rewards learned at different stages of training. Figure 4 (bottom) shows the learning curve, while Figure 4 (top) illustrates the distributions of generated rewards over the “height of the ant" feature, one dimension in the state vector.
Figure 4: Distributions of the shaped rewards over the height of the ant robot in the AntStand task at different training stages. Red diamonds represent the estimated success rate, while the blue polylines show the actual shaped rewards sampled from the Beta distribution.
Figure 5: The density of visited states in the MountainCar task for four training periods.
As learning progresses, the shaped rewards demonstrate two key attributes: the values transition from random to meaningful, and the variance decreases from uncertain to deterministic. Besides, although the sampled rewards fluctuate, their center gradually shows a positive linear correlation with the robot’s height. In early phases, the shaped rewards contain significant randomness due to higher uncertainty. Although these random signals offer limited information, they drive the agent to take tiny optimization steps in diverse directions, effectively shifting the anchors of the policies, thus broadens the actions sampled from SAC’s stochastic policy, encouraging exploration and increasing the diversity of collected samples. In later phases, the rewards converge to closely match the height of the robot, a meaningful and intuition-aligned metric, which enhances the agent’s exploitation.
To further investigate the exploration behavior of SASR, we compare the density of the visited states throughout the training in the MountainCar task with four representative exploration strategies: (1) ReLara, perturbating on both rewards and actions; (2) #Explo, rewarding novel states; (3) SAC, entropy-regularized exploration; and (4) NoisyNet (Fortunato et al.,, 2018), perturbing network weights. The density for every 25k steps is shown in Figure 5. We observe that SASR progressively covers a wider range of the state space. From 50k to 100k steps, SASR reaches positions near the goal, driven by the success rate mechanism. In contrast, ReLara covers a similar range to SASR, but is less smooth and takes longer to reach the right side. #Explo shows no clear rightward shift, as it focuses on rewarding novelty without considering the inherent value of states. SAC’s exploration is relatively narrow, making it easier to trap in local optima. NoisyNet shows a narrowing range as network weight perturbations diminish through optimization. Overall, SASR demonstrates more effective exploration and collects valuable samples sooner, leading to faster convergence.
@@ 5.3 Ablation Study
We conduct ablation studies to investigate key components of SASR. We select six representative tasks to report the experimental results, and the quantitative data is provided in Appendix D.
(a) The impact of omitting Beta distribution sampling on the performance of SASR.
(b) Learning performance of SASR with different retention rates $\phi$.
@@
(c) Learning performance of SASR with different bandwidths $h$ of Gaussian kernel.
(d) Learning performance of SASR with different RFF feature dimensions $M$.
@@
(e) Learning performance of SASR with different shaped reward weight factors $\lambda$.
Figure 6: Ablation study: the impact of key components on the performance of SASR.
Sampling from Beta distributions. (Figure 6(a)) We examine a variant of SASR that omits Beta distribution sampling, instead directly using the success rate $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$. In the early stages, limited experience makes this success rate an unreliable estimate, and using this fixed, overly confident value can mislead the agent. Furthermore, skipping Beta distribution sampling removes the exploration driven by random rewards, leading to narrower exploration. The results highlight the critical role of Beta distribution sampling for effective learning.
Retention rate $\phi$. (Figure 6(b)) The retention rate controls the estimation of counts, which affects the confidence of Beta distributions. A high retention rate ($\phi=1$) preserves all samples, resulting in a densely populated and redundant state pool, causing the Beta distribution to be prematurely overconfident, which degrades performance. Conversely, a low retention rate ($\phi=0.01$) slows convergence by requiring more iterations to gather sufficient samples. The results suggest that an appropriate retention rate is important.
Bandwidth $h$ of Gaussian kernel. (Figure 6(c)) The bandwidth $h$ controls the smoothness of the kernel functions. Beyond fixed bandwidths, we also test a linearly decreasing configuration ($h:0.5\rightarrow 0.1$), which reflects increasing confidence in KDE. Results indicate that a small bandwidth ($h=0.01$) increases the distance between samples, causing many to have zero estimated density, while a large bandwidth ($h=1$) makes samples indistinguishable due to an overly flat kernel function. Both cases result in suboptimal performance. The decreasing bandwidth setting offers no significant improvement and tends to reduce stability due to inconsistent density estimations.
@@
RFF feature dimensions $M$. (Figure 6(d)) SASR shows relatively low sensitivity to $M$, provided it is large enough to capture the states’ complexity. Results show that values like $M=500,1000,2000$ all yield similar performance levels, while significantly low dimensions, such as $M=50$, degrade performance. This is also suggested in the original RFF study (Rahimi and Recht,, 2007).
@@
Shaped reward weight factor $\lambda$. (Figure 6(e)) Regarding the shaped reward weight factor $\lambda$, we observe that SASR performs better with intermediate values like $\lambda=0.4,0.6,0.8$. At $\lambda=0.2$, the minimal shaped reward scale reduces state differentiation, leading to suboptimal performance. At $\lambda=1$, aligning the shaped reward scale with the environmental reward introduces excessive uncertainty and randomness, which may overwhelm the feedback and hinder learning. The findings highlight that maintaining a balanced reward scale is important for optimal learning outcomes.